# Benjamin S. Manning MIT John J. Horton

### v4

@@ 1 Introduction
A general, low-cost method for accurately simulating human behavior with AI agents would have wide application in the social sciences (Charness et al., 2025; Jackson et al., 2025; Anthis et al., 2025). Recognizing this potential, a growing literature explores whether large language models (LLMs) can simulate human responses in various settings.111(Argyle et al., 2022; Aher et al., 2023; Binz and Schulz, 2023; Brand et al., 2023; Park et al., 2023; Mei et al., 2024; Park et al., 2024; Chang et al., 2024; Manning et al., 2024; Capra et al., 2024; Hansen et al., 2024; Kim et al., 2024; Wang et al., 2025; Cerina and Duch, 2025; Shah et al., 2025; Broska et al., 2025; Bybee, 2025; Fish et al., 2025) Across dozens of experiments, samples of these agents respond with remarkable similarity to humans—even when simulating novel studies that did not appear in the underlying LLM’s training corpus (Hewitt et al., 2024; Binz et al., 2024; Li et al., 2024; Tranchero et al., 2024; Suh et al., 2025). Yet within this literature, others find settings where the very same models are poor human proxies.222(Santurkar et al., 2023; Atari et al., 2023; Cheng et al., 2023; Gui and Toubia, 2023; Gao et al., 2024) This inconsistency poses a challenge for AI simulations as robust and credible predictive models—particularly in settings where no prior human data exists. The core challenge is not simply achieving a close match between AI and human responses in one setting, but building agents that will generalize reliably.
@@
Despite the existence of these powerful, steerable models, constructing agents whose behavior is similar to that of real humans in a wide variety of settings is nontrivial—even when human data are available to guide the search. The set of possible prompts is vast, ranging from simple combinations of social or demographic traits to complex programmatic instructions related to how humans make decisions (Zhu et al., 2025b; Xie et al., 2025). As in other machine learning applications, the challenge is not only to avoid underfitting but also to guard against overfitting. By iterating through enough prompts, one can almost always find some arbitrary prompt that shifts the LLM’s responses to closely match a given human distribution. For example, an LLM instructed “you randomly offer between \$6 and \$9” may perfectly reproduce a distribution of human responses in a \$20 dictator game, but such a prompt would be nonsensical for a \$5 dictator game. In contrast, a prompt grounded in the underlying behavioral drivers—e.g., “you are self-interested but fair”—can perform well in-sample and plausibly extend to a range of allocation games. Standard data-driven approaches, such as a train-test split within a single dataset, cannot reliably distinguish between these two cases; the latter appears better only when tested in truly new settings. If the goal is to predict behavior in settings with no prior human data, how should researchers construct prompts?
@@
The theory-grounded agents are far better predictors of initial human play than the baseline AI off-the-shelf. In the average game, they assign $3.41$ times more probability to the actions actually taken by human subjects. The optimized agents are similarly effective compared to two well-established theoretical benchmarks: (i) a symmetric Nash equilibrium calculated for each of the 1,500 games ($2.44$ times more probability), and (ii) the game-specific predictions from the cognitive hierarchy model of Camerer et al. (2004) ($3.02$ times more probability). Because the games were randomly sampled, the corresponding confidence intervals over the relative predictive power are externally valid for the much larger population.
@@ 4.1 A pre-committed family of strategic games
The pre-committed family of games generalizes the original 11-20 money request game by parameterizing it into six independent variable components. [^19] Alsobay et al. (2025) similarly generate various public goods games and Zhu et al. (2025a) do so for over 2,000 2-by-2 strategic games. Each symmetric game preserves the core idea behind the original: two players simultaneously select a whole number between two bounds, earning guaranteed points based on their individual choice plus a potential bonus determined by both players’ choices. The six parameters—lower bound, upper bound, gap to achieve the bonus, bonus size, rule to award guaranteed points, and bonus rule—are detailed in Table 3. The table’s upper portion enumerates the possible values for the first five parameters, while the bottom section presents the eleven possible bonus rules.
@@
Upper Bound lower bound + $\{4,5,\ldots,20\}$ Max number players can select
@@ 4.3 Predictive benchmarks
Another seemingly attractive alternative would be to follow the approaches of Fudenberg and Liang (2019), Hirasawa et al. (2022), or Andrews et al. (2025), which involve training a bespoke supervised machine learning model on past game data to predict responses. However, this method is infeasible for our purposes because of two problems. First, it relies heavily on having access to a large and representative dataset, which we currently do not possess. [^22] The combined approach of Zhu et al. (2025a)—using machine learning to flexibly estimate parameters in a well-defined behavioral model and then using that model to predict responses—is infeasible for the same reason. Second, as discussed in Section 2, traditional machine learning models (e.g., regression, decision trees, or generic neural networks) are not well-suited for predicting human behavior in novel settings because they cannot adapt to structural differences across settings. Even with substantial representative data, if a game had even one new parameter value (e.g., a new bonus rule or an upper bound above the highest from the training data), the model would need to be retrained from scratch.
@@ 4.6 Results
The top panel of Figure 6 shows the estimation results. Each panel provides the histogram of the game-by-game log-likelihood ratios ($\hat{\Lambda}_{s}$) for each comparison. The vertical dashed black lines indicate the means of the log-likelihood ratios ($\bar{\Lambda}_{S}$), which are also given in the upper right corner of each panel. Bootstrapped standard errors are in parentheses. Green indicates ratios greater than zero, where the optimized AI has more predictive power, and red is the converse.
@@
Notes: The top panel shows the distribution of the log-likelihood ratios for each benchmark. The vertical black line indicates the mean, which is also provided in the upper right corner of each panel. The respective bootstrapped standard errors are in parentheses. Green indicates ratios greater than zero, where the optimized strategic AI has more predictive power, and red is the converse. The bottom panel shows the proportion of games for which the sample of strategic AI agents is the best predictor of initial play. Error bars are 95% Clopper-Pearson confidence intervals
@@
With $\varepsilon=0.2$, the strategic sample of AI agent is, on average, significantly more predictive than any of the benchmarks ($p<0.001$ for all comparisons). These differences are substantial. Starting with the leftmost panel, across all human observations in the dataset, the strategic sample of AI agents achieves an average per-observation likelihood ratio of $e^{1.23}=3.41$ in favor of the model, relative to the baseline AI. In other words, the likelihood of the observed human data under the strategic AI agent model is, on average, $3.41$ times larger per observation than under the baseline.
@@
Moving to the right, the corresponding average likelihood ratios are $e^{1.11}=3.02$ compared to the cognitive hierarchy model and $e^{0.89}=2.44$ compared to the Harsanyi-Selten-selected equilibria. To be clear, this means that the strategic AI agents outperform the cognitive hierarchy model by a larger margin than they outperform the Harsanyi-Selten Nash equilibria. This is notable because throughout the literature, the cognitive hierarchy model has been a strong predictor of initial play in strategic games. It has been explicitly shown to “explain why equilibrium theory predicts behavior well in some games and poorly in others” (Camerer et al., 2004). Of course, the cognitive hierarchy model might be more accurate with different values for $\tau$, but we had no way of knowing what these were ex ante—besides those from the literature—without the data sampled from the population we wanted to predict. Finally, the strategic AI agents show large gains compared to the random pure-strategy $e^{4.15}=63.47$ and the uniform benchmarks $e^{0.59}=1.81$.
@@
The bottom panel of Figure 6 shows the proportion of games for which the sample of strategic AI agents is the best predictor—i.e., has a positive log-likelihood ratio ($\sum_{s\in S}\mathbf{1}\{\hat{\Lambda}_{s}>0\}/|S|$). The results are consistent with the top panel. The theory-grounded sample of strategic AI agents better predicts the human responses in more games than any other model. This proportion is large and significant for the baseline ($0.715$). It is smaller, although still substantial, for both the cognitive hierarchy model ($0.64$) and the Harsanyi-Selten Nash equilibria ($0.622$).
@@ 5 Conclusion
Another exciting research direction would be to automate the theory-prediction-testing loop. Such a system would start with novel human data and relevant experimental settings, iteratively generate theory-informed candidate personas, optimize their parameters, and systematically evaluate their generalizability across samples. One can also imagine just starting with a novel setting, and then having an AI system actually search for the relevant training and validation data to execute our approach. Recent research supports the feasibility of these ideas (Xie et al., 2025; Zhu et al., 2025b). In particular, Manning et al. (2024) demonstrate AI systems capable of automating the full social scientific workflow—from hypothesis generation to experimental simulation and data analysis.
@@ D.3 Robustness checks for the family of pre-committed games in Section 4
Harsanyi-Selten Nash (….2.587*** p<.001p<.001*** p<.001p<.001*** (….0.730***
@@
Random Pure Strategy (….7.413*** p<.001p<.001*** p<.001p<.001*** (….0.942***
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.05$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Harsanyi-Selten Nash (….1.713*** p<.001p<.001*** p<.001p<.001*** (….0.679***
@@
Random Pure Strategy (….5.779*** p<.001p<.001*** p<.001p<.001*** (….0.928***
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.1$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Harsanyi-Selten Nash (….0.891*** p<.001p<.001*** p<.001p<.001*** (….0.622***
@@
Random Pure Strategy (….4.151*** p<.001p<.001*** p<.001p<.001*** (….0.902***
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.2$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Harsanyi-Selten Nash (….0.446*** p<.001p<.001*** p<.001p<.001*** (….0.575***
@@
Random Pure Strategy (….3.191*** p<.001p<.001*** p<.001p<.001*** (….0.884***
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.3$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Action Space Size 0.098∗∗∗ 0.064∗∗ -0.009 0.165∗∗∗ 0.035∗∗
@@
Action Space Size 0.088∗∗∗ 0.051∗∗ 0.001 0.150∗∗∗ 0.045∗∗∗
@@
Action Space Size 0.075∗∗∗ 0.036 0.007 0.131∗∗∗ 0.051∗∗∗
@@
Action Space Size 0.066∗∗∗ 0.024 0.008 0.115∗∗∗ 0.052∗∗∗

### v5

# arXiv:2508.17407v5[econ.GN] 01 Mar 2026

@@ 1 Introduction
A general, low-cost method for accurately simulating human behavior with AI agents would have wide application in the social sciences (Charness et al., 2025; Jackson et al., 2025; Anthis et al., 2025). Recognizing this potential, a growing literature explores whether large language models (LLMs) can simulate human responses in various settings.111(Argyle et al., 2022; Aher et al., 2023; Binz and Schulz, 2023; Brand et al., 2023; Park et al., 2023; Mei et al., 2024; Park et al., 2024; Chang et al., 2024; Manning et al., 2024; Capra et al., 2024; Hansen et al., 2024; Kim et al., 2024; Wang et al., 2025; Cerina and Duch, 2025; Shah et al., 2025; Broska et al., 2025; Bybee, 2025; Fish et al., 2025) Across dozens of experiments, samples of these agents respond with remarkable similarity to humans—even when simulating novel studies that did not appear in the underlying LLM’s training corpus (Hewitt et al., 2024; Binz et al., 2025; Li et al., 2024; Tranchero et al., 2024; Suh et al., 2025). Yet within this literature, others find settings where the very same models are poor human proxies.222(Santurkar et al., 2023; Atari et al., 2023; Cheng et al., 2023; Gui and Toubia, 2023; Gao et al., 2024) This inconsistency poses a challenge for AI simulations as robust and credible predictive models—particularly in settings where no prior human data exists. The core challenge is not simply achieving a close match between AI and human responses in one setting, but building agents that will generalize reliably.
@@
Despite the existence of these powerful, steerable models, constructing agents whose behavior is similar to that of real humans in a wide variety of settings is nontrivial—even when human data are available to guide the search. The set of possible prompts is vast, ranging from simple combinations of social or demographic traits to complex programmatic instructions related to how humans make decisions (Zhu et al., 2025a; Xie et al., 2025). As in other machine learning applications, the challenge is not only to avoid underfitting but also to guard against overfitting. By iterating through enough prompts, one can almost always find some arbitrary prompt that shifts the LLM’s responses to closely match a given human distribution. For example, an LLM instructed “you randomly offer between \$6 and \$9” may perfectly reproduce a distribution of human responses in a \$20 dictator game, but such a prompt would be nonsensical for a \$5 dictator game. In contrast, a prompt grounded in the underlying behavioral drivers—e.g., “you are self-interested but fair”—can perform well in-sample and plausibly extend to a range of allocation games. Standard data-driven approaches, such as a train-test split within a single dataset, cannot reliably distinguish between these two cases; the latter appears better only when tested in truly new settings. If the goal is to predict behavior in settings with no prior human data, how should researchers construct prompts?
@@
The theory-grounded agents are significantly better predictors of initial human play than the baseline AI off-the-shelf. In the average game, they assign $1.54$ times more probability to the actions actually taken by human subjects. The optimized agents are similarly effective compared to two well-established theoretical benchmarks: (i) a symmetric Nash equilibrium calculated for each of the 1,500 games ($1.38$ times more probability), and (ii) the game-specific predictions from the cognitive hierarchy model of Camerer et al. (2004) ($1.48$ times more probability). Because the games were randomly sampled, the corresponding confidence intervals over the relative predictive power are externally valid for the much larger population.
@@ 4.1 A pre-committed family of strategic games
The pre-committed family of games generalizes the original 11-20 money request game by parameterizing it into six independent variable components. [^19] Alsobay et al. (2025) similarly generate various public goods games and Zhu et al. (2025b) do so for over 2,000 2-by-2 strategic games. Each symmetric game preserves the core idea behind the original: two players simultaneously select a whole number between two bounds, earning guaranteed points based on their individual choice plus a potential bonus determined by both players’ choices. The six parameters—lower bound, upper bound, gap to achieve the bonus, bonus size, rule to award guaranteed points, and bonus rule—are detailed in Table 3. The table’s upper portion enumerates the possible values for the first five parameters, while the bottom section presents the eleven possible bonus rules.
@@
Upper Bound lower bound + $\{5,6,\ldots,20\}$ Max number players can select
@@ 4.3 Predictive benchmarks
Another seemingly attractive alternative would be to follow the approaches of Fudenberg and Liang (2019), Hirasawa et al. (2022), or Andrews et al. (2025), which involve training a bespoke supervised machine learning model on past game data to predict responses. However, this method is infeasible for our purposes because of two problems. First, it relies heavily on having access to a large and representative dataset, which we currently do not possess. [^22] The combined approach of Zhu et al. (2025b)—using machine learning to flexibly estimate parameters in a well-defined behavioral model and then using that model to predict responses—is infeasible for the same reason. Second, as discussed in Section 2, traditional machine learning models (e.g., regression, decision trees, or generic neural networks) are not well-suited for predicting human behavior in novel settings because they cannot adapt to structural differences across settings. Even with substantial representative data, if a game had even one new parameter value (e.g., a new bonus rule or an upper bound above the highest from the training data), the model would need to be retrained from scratch.
@@ 4.6 Results
The top panel of Figure 6 shows the estimation results. Each panel provides the histogram of the game-by-game log-likelihood ratios ($\hat{\Lambda}_{s}$) for each comparison. The vertical dashed black lines indicate the means of the log-likelihood ratios ($\bar{\Lambda}_{S}$), which are also given in the upper left corner of each panel. Bootstrapped standard errors are in parentheses. Green indicates ratios greater than zero, where the optimized AI has more predictive power, and red is the converse.
@@
Notes: The top panel shows the distribution of the log-likelihood ratios for each benchmark. The vertical black line indicates the mean, which is also provided in the upper left corner of each panel. The respective bootstrapped standard errors are in parentheses. Green indicates ratios greater than zero, where the optimized strategic AI has more predictive power, and red is the converse. The bottom panel shows the proportion of games for which the sample of strategic AI agents is the best predictor of initial play. Error bars are 95% Clopper-Pearson confidence intervals
@@
With $\varepsilon=0.2$, the strategic sample of AI agent is, on average, significantly more predictive than any of the benchmarks ($p<0.001$ for all comparisons). These differences are substantial. Starting with the leftmost panel, across all human observations in the dataset, the strategic sample of AI agents achieves an average per-observation likelihood ratio of $e^{0.43}=1.54$ in favor of the model, relative to the baseline AI. In other words, the likelihood of the observed human data under the strategic AI agent model is, on average, $1.54$ times larger per observation than under the baseline.
@@
Moving to the right, the corresponding average likelihood ratios are $e^{0.39}=1.48$ compared to the cognitive hierarchy model and $e^{0.32}=1.38$ compared to the Harsanyi-Selten-selected equilibria. To be clear, this means that the strategic AI agents outperform the cognitive hierarchy model by a larger margin than they outperform the Harsanyi-Selten Nash equilibria. This is notable because throughout the literature, the cognitive hierarchy model has been a strong predictor of initial play in strategic games. It has been explicitly shown to “explain why equilibrium theory predicts behavior well in some games and poorly in others” (Camerer et al., 2004). Of course, the cognitive hierarchy model might be more accurate with different values for $\tau$, but we had no way of knowing what these were ex ante—besides those from the literature—without the data sampled from the population we wanted to predict. Finally, the strategic AI agents show a large gain compared to the random pure-strategy benchmark $e^{1.44}=4.21$, and a more modest but still significant advantage over the uniform distribution $e^{0.2}=1.22$.
@@
The bottom panel of Figure 6 shows the proportion of games for which the sample of strategic AI agents is the best predictor—i.e., has a positive log-likelihood ratio ($\sum_{s\in S}\mathbf{1}\{\hat{\Lambda}_{s}>0\}/|S|$). The results are consistent with the top panel. The theory-grounded sample of strategic AI agents better predicts the human responses in more games than any other model. This proportion is large and significant for the baseline ($0.715$). It is smaller, although still substantial, for both the cognitive hierarchy model ($0.64$) and the Harsanyi-Selten Nash equilibria ($0.622$). Because each game has only a small number of human observations ($m_{s}$ between 1 and 5), sampling noise in $\hat{\Lambda}_{s}$ mechanically attenuates this proportion toward 0.5, almost surely making these estimates highly conservative. Empirical Bayes shrinkage methods (see Walters, 2024) could partially correct for this, but we leave the unadjusted proportions for simplicity. This issue does not affect the sample averages $\bar{\Lambda}_{S}$ reported above, which remain unbiased.
@@ 5 Conclusion
Another exciting research direction would be to automate the theory-prediction-testing loop. Such a system would start with novel human data and relevant experimental settings, iteratively generate theory-informed candidate personas, optimize their parameters, and systematically evaluate their generalizability across samples. One can also imagine just starting with a novel setting, and then having an AI system actually search for the relevant training and validation data to execute our approach. Recent research supports the feasibility of these ideas (Xie et al., 2025; Zhu et al., 2025a). In particular, Manning et al. (2024) demonstrate AI systems capable of automating the full social scientific workflow—from hypothesis generation to experimental simulation and data analysis.
@@ D.3 Robustness checks for the family of pre-committed games in Section 4
Harsanyi-Selten Nash (….0.924*** p<.001p<.001*** p<.001p<.001*** (….0.730***
@@
Random Pure Strategy (….2.574*** p<.001p<.001*** p<.001p<.001*** (….0.942***
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.05$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $\bar{\Lambda}_{S}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Harsanyi-Selten Nash (….0.614*** p<.001p<.001*** p<.001p<.001*** (….0.679***
@@
Random Pure Strategy (….2.004*** p<.001p<.001*** p<.001p<.001*** (….0.928***
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.1$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $\bar{\Lambda}_{S}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Harsanyi-Selten Nash (….0.323*** p<.001p<.001*** p<.001p<.001*** (….0.622***
@@
Random Pure Strategy (….1.436*** p<.001p<.001*** p<.001p<.001*** (….0.902***
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.2$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $\bar{\Lambda}_{S}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Harsanyi-Selten Nash (….0.165*** p<.001p<.001*** p<.001p<.001*** (….0.575***
@@
Random Pure Strategy (….1.102*** p<.001p<.001*** p<.001p<.001*** (….0.884***
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.3$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $\bar{\Lambda}_{S}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Action Space Size 0.030∗∗∗ 0.020 -0.004 0.049∗∗∗ 0.011
@@
Action Space Size 0.027∗∗∗ 0.016 -0.001 0.046∗∗∗ 0.015∗∗∗
@@
Action Space Size 0.024∗∗∗ 0.012 0.002 0.041∗∗∗ 0.017∗∗∗
@@
Action Space Size 0.021∗∗∗ 0.008 0.002 0.037∗∗∗ 0.017∗∗∗